# Adaptive Selection for Homogeneous Tools: An Instantiation in the RAG Scenario

### v2

@@ Appendix B Data Processing Details and Statistics
FreshQA will be regularly updated by its creator. We have used the version updated at 2024-05-20. We use the first 2000 examples in the official HotpotQA testset, and the whole FreshQA datasets as test cases. We randomly select 20000 examples from the official HotpotQA train set as training cases. For WebQA and CDQA, we use the official data split. For each example, we first use different search tools and LLM for retrieval-augmented generation. We discard examples with invalid search results. After obtaining labels, we merge training examples of CDQA and WebQA, and randomly split the Train/Dev partitions with a ratio of 0.85/0.15. We train the model on the mixed train part and test our model on the testsets of WebQA, TimeQA, and CDQA. Similarly, we train our model on HotpotQA and test it on the testsets of HotpotQA and FreshQA. The statistics of used datasets are shown in Table 3.
@@
For the English experiment, we adopt FacebookAI/roberta-base as the backbone. For the Chinese experiment, we adopt damo/nlp_roberta_backbone_base_std as the backbone. These pre-trained models can be downloaded from the huggingface model-hup. We use the Adam optimizer and linearly decrease the learning rate to zero with a 10% warmup ratio. We use PyTorch toolkit to conduct all experiments on the Ubuntu server with a V100 (32G) GPU. All the hyperparameters for are searched according to the final QA accuracy on the development set. For reproduction, we set the random seed to 42 for all experiments. The searched parameters for two datasets are shown in Table 4.
English Experiment Chinese Experiment
batch size 16 16
num-epochs 20 20
$lr$ 5e-5 5e-5
$\alpha$ 0.5 0.8
$k$ 2 2
Table 4: The used hyperparameters for English and Chinese experiments.
Appendix C More Cost-accuracy Curves
(a) The cost-accuracy curve on the CDQA test.
(b) The cost-accuracy curve on the TimeQA test.
Figure 5: The cost-accuracy curves on the testsets of TimeQA and CDQA.
We additionally present the cost-accuracy curves on the testsets of CDQA and TimeQA. The result is shown in Figure 5. We have the findings that coincide with Section 4.3.

### v3

@@ Appendix A Data Processing Details and Statistics
For WebQA and CDQA, we use the official data split. For each example, we first use different search tools and LLMs for retrieval-augmented generation. We discard examples with invalid search results. After obtaining labels, we merge training examples of CDQA and WebQA, and randomly split the Train/Dev partitions with a ratio of 0.85/0.15. We train the model on the mixed train part and test our model on the testsets of WebQA, PrivateTimeQA, and CDQA. The statistics of used datasets are shown in Table 3.
@@
We adopt damo/nlp_roberta_backbone_large_std as the backbone. These pre-trained models can be downloaded from the modelscope [^4] https://www.modelscope.cn/home. We use the Adam optimizer and linearly decrease the learning rate to zero with a 10% warmup ratio. We use PyTorch toolkit to conduct all experiments on the Ubuntu server with a V100 (32G) GPU. All the hyperparameters for are searched according to the final QA accuracy on the development set. For reproduction, we set the random seed to 42 for all experiments. The searched parameters are shown in Table 4.
batch size 16
num-epochs 20
$lr$ 5e-5
$\alpha$ 0.8
$k$ 2
Table 4: The used hyperparameters in our experiments.